An alloy with density 2.85 g/cm³ consists of 93.7% Al and 6.3%Cu by mass. 0.691cm³ of alloy reacts with HCl(aq). If all Al and none Cu reacts with acid, find mass of H2 obtained.
1
Expert's answer
2020-12-10T07:45:05-0500

Solution.

"m = \rho \times V"

m = 1.97 g

"m(Al) = 0.937 \times 1.97 = 1.85 g"

"n(H2) = \frac{n(Al) \times 3}{2}"

"n(Al) = \frac{1.85}{26.98} = 0.07 \ mol"

n(H2) = 0.105 mol

"m(H2) = 0.105 \times 2 = 0.21 g"

Answer:

m(H2) = 0.21 g